OpenStudy (anonymous):

I have two circles. One has a radius of 2 and the other, 4. How can I find area and figure out their equations? BTW, the center for both is at (6,4)

OpenStudy (lgbasallote):

area = \(\pi r^2\)

OpenStudy (anonymous):

Hello, Facepalm.:) :P

OpenStudy (lgbasallote):

what's the problem?

OpenStudy (anonymous):

Thanks! :D I need to know the formula for finding the equation of the circles. Their center at (6,4).

OpenStudy (lgbasallote):

oh lol...you should memorize those formulas...they are very common in later maths

OpenStudy (anonymous):

Yeah. :|

OpenStudy (anonymous):

I meant the equation of the circle itself. :S It's like, (x + 5)(y + 5) = r or something.

OpenStudy (lgbasallote):

oh lol \( (x-h)^2 + (y - k)^2 = r^2\) were (h,k) = center

OpenStudy (lgbasallote):

that's the distance formula actually :p

OpenStudy (lgbasallote):

 see?

OpenStudy (lgbasallote):

just some additional info...not really essential

OpenStudy (anonymous):

OH! Thank you! :DDDDD